media%2F1b0%2F1b08b91b-07b4-4f0a-bc39-4a
Answer #1

The x and y componets of velocity of plane are,

u_x= u \times \cos \theta= 99 \times \cos 22=91.79 ms^{-1}

u_y= u \times \sin \theta= 99 \times \sin 22= 37.086ms^{-1}

Consider the motion of the bag along y axis,

H= u_yt+\frac{1}{2}(-g)t^2

-109= 37.086t-\frac{1}{2}9.8t^2

Solve for t,

using t=\frac{ -b \pm \sqrt{b^2-4ac}}{2a}

where a=4.9, b=-37.086 and c=-109

take the positive value of t, which is 9.83s

Now consider the horizontal motion of suit case,

There is no acceleration during horizontal motion,

Therefore s= u_x \times t

s= 91.79 \times 9.83=902.3 m

The suit case falls at a distance 902.3 m from dog.

Note:

The general sign convention says that all upward motion are considered to be positive and all downward motions are considered to be negative.

So, in this problem the suit case moves downwards, so H=-109m and g=-9.8ms^{-2}